• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name ulmer

    Log entry time 11:59:37 on June 1,2002

    Entry number 4350

    This entry is a followup to: 4317

    keyword=ARC energy/Hall C tgt cool

    The crew chief says that AES (now EES) has been notified, as well as "software on call" for the problem with the ARC energy measurement. I verified with Pavel (the PD) that people will look into this problem and told him that it's essential we get this fixed by Monday, since we would like to do the energy measurement then (see below).

    Related to scheduling this energy measurement is the Hall C target cooldown, scheduled for Monday. This may produce interruptions to our target cooling during two 1 hour periods, spaced by 3 hours. We have been asked to specify the time of the first such period which is convenient for us - i.e. we would be best served by doing non-cryotarget related stuff during this time. Pavel tells me that we can take beam on other targets or (related to above) we could do the ARC energy measurement during this time. This obviously hinges on our having the above mentioned problem fixed.